OpenStudy (lisa123):

Write equations of the lines through the given point parallel to and perpendicular to the given line. 6x − 2y = 9, (3, 1)

OpenStudy (anonymous):

ok find a slope from this line and since the parallel slope has identical slope , u can use that to get an equation by using

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

do u know hoe to find a slope from an equation , you make y the subject like (y=...)

OpenStudy (anonymous):

how*

OpenStudy (lisa123):

I think I did this but I got it wrong 6x-2y=9 -2y=-6x+9 y=3x-9/2 y-1=3(x-3) y=3x-8:parallel line

OpenStudy (anonymous):

well done u are right :)

OpenStudy (lisa123):

No, the answer is wrong...according to my hw website

OpenStudy (anonymous):

yea i realized u didnt do the third step right

OpenStudy (lisa123):

oh what is that?

OpenStudy (anonymous):

no ur right the book is wrong

OpenStudy (anonymous):

i thought u maybe messed up while making the y subject its right

OpenStudy (lisa123):

i also got y=-1/3x for the perpendicular line but it tells me thats wrong too

OpenStudy (anonymous):

nope ur right

OpenStudy (lisa123):

Hmm okay I'll show my teacher. Thanks!

OpenStudy (anonymous):

i am 110% ur right

OpenStudy (anonymous):

no worries happy to help

OpenStudy (lisa123):

Can you help me with another problem

OpenStudy (anonymous):

yea sure

OpenStudy (lisa123):

Find the x- and y-intercepts of the graph of the equation. y = x4 − 25

OpenStudy (lisa123):

x - intercept( larger value): sqrt 5,0 y-intercept : 0,-25 x- intercept (smaller value) ??

OpenStudy (anonymous):

is the eq \[x ^{4}-25\]